45. Chapter ma2pe08r, section.271, Problem 208 (ID: 208.27.1- MC-MANK08) If the annual interest rate is 2%, what is the present value of the savings bond? Suppose you own a savings bond that will pay you $100 in 7 years. O a. $27.91 b. $87.06 O c. $93.64 O d. $87.06
Answer #1

The present value is $87.06

The formula used is :-

A = P(1 + rt)

Here , A = total amount payable after time period

P = pricipal amount

r = Rate of interest

t = time period

Hence in question A = 100 , P = ? , r = 2/100 , t = 7 years

Using values in formula,

100 = P ( 1 + 2/100 ×7 )

100 = P ( 1 + 14/100 )

100 = P ( 114/100 )

100 ×100 /114 = P

10000/ 114 = P

Therefore P = 87.7 (approx)